OpenStudy (anonymous):

Write a system of equations for each problem. Kami has $2.35 in nickels and dimes. If she has 33 coins in all, find the number of nickels and dimes.

OpenStudy (anonymous):

N = nickels D = dimes N+D=33 (0.05)N+(0.1)D=2.35

OpenStudy (anonymous):

solve for N and D if you know how to, and then N and D will be the number of nickels and dimes respectively